OpenStudy (anonymous):

trying to help son, and totally forgot how to do all of this. what are the steps in (a^5a^9)^-5/a^-15

OpenStudy (anonymous):

(a^(5+9))^-5 /a^-15 ===> a^(14*(-5)) / a^-15 ===> a^-70/a^-15===> a^(-70+15) = a^(-55) = 1/a^(55)

OpenStudy (anonymous):

Thanks I get it , I guess my son had written down the wrong answer so I kept trying to get the equation to result in a^-4. This really helped

OpenStudy (anonymous):

yw